OpenStudy (anonymous):

find y' if 3xy^2+y^3=e^(x+y)

OpenStudy (anonymous):

derivative on both sides: (3xy^2)' = 3y^2 + 6xyy', (y^3)' = 3y^2*y' , (6xy)' = 6y + 6xy', (e^(x+y))' = e^(x+y) * (1 + y') Then, 3y^2 + 6xyy' + 3y^2y' = e^(x+y)*(1+y') Then, \[y' = {(3y^2 - e ^{x+y})} / ( {e^{x+y} - 3y^2 - 6xy})\]